OpenStudy (anonymous):

f(x)=x(x^2-x-2) Determine whether the Mean Value Theorem can be applied to f on the closed interval [a,b]. If the Mean Value Theorem can be applied, find all values of c in the open interval (a,b) such that f'(c)=f(b)-f(a)/b-a

OpenStudy (anonymous):

Check if the hypothesis of the MVT is satisfied by f. Then just solve the equation for c

OpenStudy (anonymous):

is it \[f(x)=x(x^2-x-2)=x^3-x^2-2x\]? all polynomials are continuous for all real numbers, so the mvt applies to any polynomial over any interval

OpenStudy (anonymous):

do you have specific values for \(a\) and \(b\) ?

OpenStudy (anonymous):

The interval is [-1,1]

OpenStudy (anonymous):

is it the function i wrote?

OpenStudy (anonymous):

Yes, that is the function

OpenStudy (anonymous):

the cubic function ok lets take it step by step

OpenStudy (anonymous):

Yes, please!

OpenStudy (anonymous):

you need \(f(1)=1^3-1^2-2\times 1=-2\)

OpenStudy (anonymous):

I wish I understood that. This is what I've done: f'(x)= 3x^2-2x-2 f(-1)=3(-1)-2(-1)-2=3 f(1)=3(1)^2-2(1)-2=-1

OpenStudy (anonymous):

Oh okay, I see what you did

OpenStudy (anonymous):

you also need \(f(-1)=(-1)^3-(-1)^2-2\times (-1)=0\)

OpenStudy (anonymous):

it is the function evaluated at the endpoint of the derivative, not the derivative

OpenStudy (anonymous):

so \(f(b)-f(a)=f(1)-f(-1)=-2-0=-2\)

OpenStudy (anonymous):

and \[\frac{f(b)-f(a)}{b-a}=\frac{f(1)-f(-1)}{1-(-1)}=\frac{-2}{2}=-1\]

OpenStudy (anonymous):

your final job is to take the derivative, set it equal to \(-1\) and solve the resulting (quadratic) equation you good from there?

OpenStudy (anonymous):

So 3x^2-2x-2=-1 will I solve this using quadratic equation? (so sorry you're having to simplify this so much!)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the derivative is \(f'(x)=3x^2-2x-2\) and you want to know for what number in the interval \([-1,1]\) that it is equal to \(-1\) so set \[3x^2-2x-2=-1\] and solve for \(x\) via \[3x^2-2x-1=0\] etc

OpenStudy (anonymous):

by some miracle this problem was cooked up to factor, so you can use \[(3x+1)(x-1)=0\] to solve it

OpenStudy (anonymous):

Thank you so, SO much!